# A Population Analysis of 20 Exoplanets Observed from the Optical to the Near-infrared Wavelengths with HST: Evidence for Widespread Stellar Contamination

### v2

@@ Appendix E Retrieval results
| H2O (base model) | 247.310.21 | |||
| H2O + clouds | 252.340.18 | |||
| H2O + Stellar activity | 240.760.30 | |||
| H2O + Stellar activity + clouds | 243.900.29 | |||
@@
| H2O (base model) | 240.740.21 | |||
| H2O + clouds | 246.360.18 | |||
| H2O + Stellar activity | 239.590.31 | |||
| H2O + Stellar activity + clouds | 243.570.29 | |||
@@
| H2O (base model) (sol 0) | 0.34650.0011 | 243.680.23 | ||
| H2O (base model) (sol 1) | 247.570.23 | |||
| H2O + clouds | 255.590.20 | |||
| H2O + clouds + Na + K | 257.120.21 | |||
| H2O + Stellar activity | 249.460.24 | |||
| H2O + Stellar activity + clouds | 256.800.21 | |||
| H2O + Stellar activity + clouds + Na + K | 256.770.21 | |||
@@
| H2O (base model) (sol 0) | 252.540.23 | |||
| H2O (base model) (sol 1) | 241.910.23 | |||
| H2O + clouds | 260.360.20 | |||
| H2O + clouds + Na + K | 259.360.20 | |||
| H2O + Stellar activity | 253.770.25 | |||
| H2O + Stellar activity + clouds | 259.300.21 | |||
| H2O + Stellar activity + clouds + Na + K | 258.560.22 | |||
@@
| H2O (base model) (sol 0) | 250.620.23 | |||
| H2O (base model) (sol 1) | 241.340.23 | |||
| H2O + clouds | 258.100.20 | |||
| H2O + clouds + Na + K | 257.290.20 | |||
| H2O + Stellar activity | 248.580.25 | |||
| H2O + Stellar activity + clouds | 255.340.21 | |||
| H2O + Stellar activity + clouds + Na + K | 255.250.21 | |||
@@
| H2O (base model) | 269.740.21 | |||
| H2O + Na (sol 0) | 282.850.23 | |||
| H2O + clouds | 272.790.23 | |||
| H2O + clouds + Na (sol 0) | 283.100.24 | |||
| H2O + clouds + Na (sol 1) | 286.400.23 | |||
| H2O + Stellar activity (spots+faculae) | 295.150.25 | |||
| H2O + Stellar activity (spots only) | 288.170.24 | |||
| H2O + Stellar activity (spots+faculae) + Na | 1.18980.0157 | 294.750.25 | ||
| H2O + Stellar activity (spots only) + Na | 288.050.24 | |||
| H2O + Stellar activity (spots+faculae) + clouds | 295.830.25 | |||
| H2O + Stellar activity (spots only) + clouds | 293.990.24 | |||
| H2O + Stellar activity (spots+faculae) + clouds + Na | 295.150.25 | |||
| H2O + Stellar activity (spots only) + clouds + Na | 293.630.24 | |||
@@
| H2O (base model) | 272.860.21 | |||
| H2O + Na | 289.330.23 | |||
| H2O + clouds | 272.080.21 | |||
| H2O + clouds + Na | 289.200.23 | |||
| H2O + Stellar activity (spots+faculae) | 292.750.25 | |||
| H2O + Stellar activity (spots only) | 282.560.24 | |||
| H2O + Stellar activity (spots+faculae) + Na | 292.390.25 | |||
| H2O + Stellar activity (spots only) + Na | 288.120.24 | |||
| H2O + Stellar activity (spots+faculae) + clouds | 293.000.25 | |||
| H2O + Stellar activity (spots only) + clouds | 291.330.25 | |||
| H2O + Stellar activity (spots+faculae) + clouds + Na | 293.930.24 | |||
| H2O + Stellar activity (spots only) + clouds + Na | 291.780.25 | |||
@@
| H2O (base model) | 278.650.21 | |||
| H2O + clouds | 278.850.20 | |||
| H2O + Stellar activity | 275.100.23 | |||
| H2O + Stellar activity + clouds | 275.530.22 | |||
@@
| H2O (base model) | 274.830.21 | |||
| H2O + clouds | 276.140.19 | |||
| H2O + Stellar activity | 274.590.23 | |||
| H2O + Stellar activity + clouds | 274.990.21 | |||
@@
| H2O (base model) | 264.090.21 | |||
| H2O + clouds | 264.410.21 | |||
| H2O + clouds + Na | 274.910.22 | |||
| H2O + Stellar activity | 274.750.24 | |||
| H2O + Stellar activity + clouds | 278.730.22 | |||
| H2O + Stellar activity + clouds + Na | 278.830.23 | |||
@@
| H2O (base model) (sol 0) | 401.070.20 | |||
| H2O (base model) (sol 1) | 400.320.20 | |||
| H2O + clouds | 403.410.21 | |||
| H2O + clouds + K + Na | 402.730.22 | |||
| H2O + Stellar activity | 401.160.22 | |||
| H2O + Stellar activity + clouds | 402.890.22 | |||
| H2O + Stellar activity + clouds + K + Na | 401.510.23 | |||
@@
| H2O (base model) | 0.51990.0029 | 397.980.21 | ||
| H2O + clouds | 398.250.22 | |||
| H2O + clouds + K + Na | 396.790.23 | |||
| H2O + Stellar activity | 395.930.23 | |||
| H2O + Stellar activity + clouds | 398.610.23 | |||
| H2O + Stellar activity + clouds + K + Na | 397.520.23 | |||
@@
| H2O (base model) | 314.070.23 | |||
| H2O + clouds | 315.710.23 | |||
| H2O + Stellar activity | 314.460.23 | |||
| H2O + Stellar activity + clouds | 0.25 | 316.140.25 | ||
@@
| H2O (base model) | 0.17 | 318.310.23 | ||
| H2O + clouds (sol0) | 319.650.23 | |||
| H2O + clouds (sol1) | 310.970.27 | |||
| H2O + Stellar activity | 319.370.22 | |||
| H2O + Stellar activity + clouds | 319.440.24 | |||
@@
| H2O (base model) | 291.840.22 | |||
| H2O + clouds | 291.160.22 | |||
| H2O + Na | 303.330.25 | |||
| H2O + clouds + Na | 307.960.27 | |||
| H2O + Stellar activity | 291.840.22 | |||
| H2O + Stellar activity + clouds | 291.150.23 | |||
| H2O + Stellar activity + Na | 305.950.24 | |||
| H2O + Stellar activity + clouds + Na | 306.760.25 | |||
@@
| H2O (base model) | 294.420.22 | |||
| H2O + clouds | 295.690.22 | |||
| H2O + Na | 310.960.25 | |||
| H2O + clouds + Na | 324.870.25 | |||
| H2O + Stellar activity | 295.200.22 | |||
| H2O + Stellar activity + clouds | 296.460.22 | |||
| H2O + Stellar activity + Na | 309.940.26 | |||
| H2O + Stellar activity + clouds + Na | 325.610.24 | |||
@@
| H2O (base model) | 367.450.22 | |||
| H2O + clouds (sol 0) | 375.710.23 | |||
| H2O + clouds (sol 1) | 371.890.23 | |||
| H2O + clouds + K (sol 0) | 371.570.23 | |||
| H2O + clouds + K (sol 1) | 375.050.23 | |||
| H2O + Stellar activity | 374.180.24 | |||
| H2O + Stellar activity + clouds | 377.260.24 | |||
| H2O + Stellar activity + clouds + K | 376.860.24 | |||
@@
| H2O (base model) | 380.370.25 | |||
| H2O + clouds | 383.110.25 | |||
| H2O + clouds + K | 384.230.25 | |||
| H2O + Stellar activity | 386.710.24 | |||
| H2O + Stellar activity + clouds | 389.740.25 | |||
| H2O + Stellar activity + clouds + K | 391.350.25 | |||
@@
| H2O (base model) | 306.930.19 | |||
| H2O + clouds | 307.140.18 | |||
| H2O + clouds + Na (sol 0) | 307.480.18 | |||
| H2O + clouds + Na (sol 1) | 305.300.18 | |||
| H2O + Stellar activity | 305.220.21 | |||
| H2O + Stellar activity + clouds | 306.360.20 | |||
| H2O + Stellar activity + clouds + Na | 306.480.20 | |||
@@
| H2O (base model) | 300.530.19 | |||
| H2O + clouds | 301.030.18 | |||
| H2O + clouds + Na (sol 0) | 300.670.17 | |||
| H2O + clouds + Na (sol 1) | 299.910.17 | |||
| H2O + Stellar activity | 297.490.22 | |||
| H2O + Stellar activity + clouds | 298.750.20 | |||
| H2O + Stellar activity + clouds + Na | 298.700.21 | |||
@@
| H2O (base model) | 273.620.19 | |||
| H2O + clouds | 275.000.18 | |||
| H2O + Stellar activity | 271.840.21 | |||
| H2O + Stellar activity + clouds | 272.690.20 | |||
@@
| H2O (base model) (sol 0) | 270.500.20 | |||
| H2O (base model) (sol 1) | 268.050.22 | |||
| H2O + clouds (sol 0) | 272.900.20 | |||
| H2O + clouds (sol 1) | 269.110.21 | |||
| H2O + Stellar activity | 269.560.21 | |||
| H2O + Stellar activity + clouds | 271.600.22 | |||
@@
| H2O (base model) | 239.380.22 | |||
| H2O + clouds (sol 0) | 227.490.21 | |||
| H2O + clouds (sol 1) | 227.380.21 | |||
| H2O + Stellar activity | 266.220.24 | |||
| H2O + Stellar activity + clouds | 265.600.25 | |||
@@
| H2O (base model) | 242.570.22 | |||
| H2O + clouds | 241.960.23 | |||
| H2O + Stellar activity | 263.830.25 | |||
| H2O + Stellar activity + clouds | 263.180.25 | |||
@@
| H2O (base model) | 311.710.22 | |||
| H2O + clouds (sol 0) | 312.630.23 | |||
| H2O + clouds (sol 1) | 298.600.23 | |||
| H2O + Stellar activity | 309.700.25 | |||
| H2O + Stellar activity + clouds | 314.100.25 | |||
@@
| H2O (base model) | 286.900.24 | |||
| H2O + clouds (sol 0) | 289.170.25 | |||
| H2O + clouds (sol 1) | 288.640.23 | |||
| H2O + Stellar activity | 289.680.26 | |||
| H2O + Stellar activity + clouds | 313.790.25 | |||
@@
| H2O (base model) | 304.500.19 | |||
| H2O + clouds (sol 0) | 305.860.20 | |||
| H2O + clouds (sol 1) | 302.880.20 | |||
| H2O + clouds + K (sol 0) | 306.470.20 | |||
| H2O + clouds + K (sol 1) | 303.260.21 | |||
| H2O + Stellar activity | 311.810.21 | |||
| H2O + Stellar activity + clouds | 312.400.20 | |||
| H2O + Stellar activity + clouds + K | 312.590.21 | |||
@@
| H2O (base model) | 302.890.19 | |||
| H2O + clouds (sol 0) | 302.600.18 | |||
| H2O + clouds (sol 1) | 302.080.19 | |||
| H2O + clouds + K (sol 0) | 304.020.19 | |||
| H2O + clouds + K (sol 1) | 302.600.19 | |||
| H2O + Stellar activity | 303.510.22 | |||
| H2O + Stellar activity + clouds | 304.110.21 | |||
| H2O + Stellar activity + clouds + K | 305.190.21 | |||
@@
| H2O (base model) | 289.580.22 | |||
| H2O + clouds | 291.030.21 | |||
| H2O + Stellar activity | 288.710.25 | |||
| H2O + Stellar activity + clouds | 289.730.24 | |||
@@
| H2O (base model) | 288.140.23 | |||
| H2O + clouds (sol 0) | 288.560.23 | |||
| H2O + clouds (sol 1) | 285.550.22 | |||
| H2O + Stellar activity | 284.940.26 | |||
| H2O + Stellar activity + clouds | 286.270.26 | |||
@@
| H2O (base model) | 286.160.20 | |||
| H2O + clouds (sol 0) | 293.050.21 | |||
| H2O + clouds (sol 1) | 291.120.21 | |||
| H2O + clouds + K (sol 0) | 292.860.22 | |||
| H2O + clouds + K (sol 1) | 290.750.22 | |||
| H2O + Stellar activity | 295.940.22 | |||
| H2O + Stellar activity + clouds | 298.550.20 | |||
| H2O + Stellar activity + clouds + K | 298.260.20 | |||
@@
| H2O (base model) | 303.000.19 | |||
| H2O + clouds (sol 0) | 304.220.18 | |||
| H2O + clouds (sol 1) | 302.090.19 | |||
| H2O + clouds + K | 303.880.19 | |||
| H2O + Stellar activity | 305.740.22 | |||
| H2O + Stellar activity + clouds | 308.170.19 | |||
| H2O + Stellar activity + clouds + K | 308.030.20 | |||
@@
| H2O (base model) | 293.620.22 | |||
| H2O + clouds (sol 0) | 303.640.23 | |||
| H2O + clouds (sol 1) | 307.450.22 | |||
| H2O + Stellar activity (sol 0) | 297.240.24 | |||
| H2O + Stellar activity (sol 1) | 306.340.30 | |||
| H2O + Stellar activity + clouds | 307.580.26 | |||
@@
| H2O (base model) (sol 0) | 302.790.21 | |||
| H2O (base model) (sol 1) | 306.540.22 | |||
| H2O + clouds | 314.240.21 | |||
| H2O + Stellar activity (sol 0) | 312.950.29 | |||
| H2O + Stellar activity (sol 1) | 304.920.25 | |||
| H2O + Stellar activity + clouds | 314.160.25 | |||
@@
| H2O (base model) (sol 0) | 255.670.22 | |||
| H2O (base model) (sol 1) | 473.608.26 | 253.420.23 | ||
| H2O + clouds | 268.570.24 | |||
| H2O + clouds + K + Na | 273.700.26 | |||
| H2O + Stellar activity | 255.220.25 | |||
| H2O + Stellar activity + clouds | 280.940.27 | |||
| H2O + Stellar activity + clouds + Na + K | 281.450.27 | |||

### v3

@@ Appendix E Retrieval results
H2O (base model) | 247.30.2 | |||
H2O + clouds | 252.30.2 | |||
H2O + Stellar activity | 240.80.3 | |||
H2O + Stellar activity + clouds | 243.90.3 | |||
@@
H2O (base model) | 240.70.2 | |||
H2O + clouds | 246.40.2 | |||
H2O + Stellar activity | 239.60.3 | |||
H2O + Stellar activity + clouds | 243.60.3 |
@@
H2O (base model) (sol 0) | 243.70.2 | |||
H2O (base model) (sol 1) | 247.60.2 | |||
H2O + clouds | 255.590.2 | |||
H2O + clouds + Na + K | 257.120.2 | |||
H2O + Stellar activity | 249.460.2 | |||
H2O + Stellar activity + clouds | 256.80.2 | |||
H2O + Stellar activity + clouds + Na + K | 256.770.2 | |||
@@
H2O (base model) (sol 0) | 252.50.2 | |||
H2O (base model) (sol 1) | 241.90.2 | |||
H2O + clouds | 260.40.2 | |||
H2O + clouds + Na + K | 259.40.2 | |||
H2O + Stellar activity | 253.80.3 | |||
H2O + Stellar activity + clouds | 259.30.2 | |||
H2O + Stellar activity + clouds + Na + K | 258.60.2 | |||
@@
H2O (base model) (sol 0) | 250.60.2 | |||
H2O (base model) (sol 1) | 241.30.2 | |||
H2O + clouds | 258.10.2 | |||
H2O + clouds + Na + K | 257.30.2 | |||
H2O + Stellar activity | 248.60.3 | |||
H2O + Stellar activity + clouds | 255.30.2 | |||
H2O + Stellar activity + clouds + Na + K | 255.30.2 |
@@
H2O (base model) | 269.70.2 | |||
H2O + Na (sol 0) | 282.90.2 | |||
H2O + clouds | 272.80.2 | |||
H2O + clouds + Na (sol 0) | 283.10.2 | |||
H2O + clouds + Na (sol 1) | 286.40.2 | |||
H2O + Stellar activity (spots+faculae) | 295.20.3 | |||
H2O + Stellar activity (spots only) | 288.20.2 | |||
H2O + Stellar activity (spots+faculae) + Na | 294.80.3 | |||
H2O + Stellar activity (spots only) + Na | 288.10.2 | |||
H2O + Stellar activity (spots+faculae) + clouds | 295.80.3 | |||
H2O + Stellar activity (spots only) + clouds | 294.00.2 | |||
H2O + Stellar activity (spots+faculae) + clouds + Na | 295.20.3 | |||
H2O + Stellar activity (spots only) + clouds + Na | 293.60.2 | |||
@@
H2O (base model) | 272.90.2 | |||
H2O + Na | 289.30.2 | |||
H2O + clouds | 272.10.2 | |||
H2O + clouds + Na | 289.20.2 | |||
H2O + Stellar activity (spots+faculae) | 292.80.3 | |||
H2O + Stellar activity (spots only) | 282.60.2 | |||
H2O + Stellar activity (spots+faculae) + Na | 292.40.3 | |||
H2O + Stellar activity (spots only) + Na | 288.10.2 | |||
H2O + Stellar activity (spots+faculae) + clouds | 293.00.3 | |||
H2O + Stellar activity (spots only) + clouds | 291.30.3 | |||
H2O + Stellar activity (spots+faculae) + clouds + Na | 293.90.2 | |||
H2O + Stellar activity (spots only) + clouds + Na | 291.80.3 |
@@
H2O (base model) | 278.70.2 | |||
H2O + clouds | 278.90.2 | |||
H2O + Stellar activity | 275.10.2 | |||
H2O + Stellar activity + clouds | 275.50.2 | |||
@@
H2O (base model) | 274.80.2 | |||
H2O + clouds | 276.10.2 | |||
H2O + Stellar activity | 274.60.2 | |||
H2O + Stellar activity + clouds | 2750.2 |
@@
H2O (base model) | 264.10.2 | |||
H2O + clouds | 264.40.2 | |||
H2O + clouds + Na | 274.90.2 | |||
H2O + Stellar activity | 274.80.2 | |||
H2O + Stellar activity + clouds | 278.70.2 | |||
H2O + Stellar activity + clouds + Na | 278.80.2 |
@@
H2O (base model) (sol 0) | 401.10.2 | |||
H2O (base model) (sol 1) | 400.30.2 | |||
H2O + clouds | 403.40.2 | |||
H2O + clouds + K + Na | 402.70.2 | |||
H2O + Stellar activity | 401.20.2 | |||
H2O + Stellar activity + clouds | 402.90.2 | |||
H2O + Stellar activity + clouds + K + Na | 401.50.2 | |||
@@
H2O (base model) | 0.5200.003 | 398.00.2 | ||
H2O + clouds | 398.30.2 | |||
H2O + clouds + K + Na | 396.80.2 | |||
H2O + Stellar activity | 395.90.2 | |||
H2O + Stellar activity + clouds | 398.60.2 | |||
H2O + Stellar activity + clouds + K + Na | 397.50.2 |
@@
H2O (base model) | 314.10.2 | |||
H2O + clouds | 315.70.2 | |||
H2O + Stellar activity | 314.50.2 | |||
H2O + Stellar activity + clouds | 316.10.3 | |||
@@
H2O (base model) | 318.30.2 | |||
H2O + clouds (sol0) | 319.70.2 | |||
H2O + clouds (sol1) | 311.00.3 | |||
H2O + Stellar activity | 319.40.2 | |||
H2O + Stellar activity + clouds | 319.40.2 | |||
@@
H2O (base model) | 291.80.2 | |||
H2O + clouds | 291.20.2 | |||
H2O + Na | 303.30.3 | |||
H2O + clouds + Na | 308.00.3 | |||
H2O + Stellar activity | 291.80.2 | |||
H2O + Stellar activity + clouds | 291.20.2 | |||
H2O + Stellar activity + Na | 306.00.2 | |||
H2O + Stellar activity + clouds + Na | 306.80.3 | |||
@@
H2O (base model) | 294.40.2 | |||
H2O + clouds | 2960.2 | |||
H2O + Na | 3110.3 | |||
H2O + clouds + Na | 324.90.3 | |||
H2O + Stellar activity | 295.20.2 | |||
H2O + Stellar activity + clouds | 296.50.2 | |||
H2O + Stellar activity + Na | 309.90.3 | |||
H2O + Stellar activity + clouds + Na | 325.60.2 |
@@
H2O (base model) | 367.50.2 | |||
H2O + clouds (sol 0) | 375.70.2 | |||
H2O + clouds (sol 1) | 371.90.2 | |||
H2O + clouds + K (sol 0) | 371.60.2 | |||
H2O + clouds + K (sol 1) | 375.10.2 | |||
H2O + Stellar activity | 374.20.2 | |||
H2O + Stellar activity + clouds | 377.30.2 | |||
H2O + Stellar activity + clouds + K | 376.90.2 | |||
@@
H2O (base model) | 380.40.3 | |||
H2O + clouds | 383.10.3 | |||
H2O + clouds + K | 384.20.3 | |||
H2O + Stellar activity | 386.70.2 | |||
H2O + Stellar activity + clouds | 389.70.3 | |||
H2O + Stellar activity + clouds + K | 391.40.3 |
@@
H2O (base model) | 306.90.2 | |||
H2O + clouds | 307.10.2 | |||
H2O + clouds + Na (sol 0) | 307.50.2 | |||
H2O + clouds + Na (sol 1) | 305.30.2 | |||
H2O + Stellar activity | 305.20.2 | |||
H2O + Stellar activity + clouds | 306.40.2 | |||
H2O + Stellar activity + clouds + Na | 306.50.2 | |||
@@
H2O (base model) | 300.50.2 | |||
H2O + clouds | 301.00.2 | |||
H2O + clouds + Na (sol 0) | 300.70.2 | |||
H2O + clouds + Na (sol 1) | 299.90.2 | |||
H2O + Stellar activity | 297.50.2 | |||
H2O + Stellar activity + clouds | 298.80.2 | |||
H2O + Stellar activity + clouds + Na | 298.70.2 |
@@
H2O (base model) | 273.60.2 | |||
H2O + clouds | 275.00.2 | |||
H2O + Stellar activity | 271.80.2 | |||
H2O + Stellar activity + clouds | 272.70.2 | |||
@@
H2O (base model) (sol 0) | 270.50.2 | |||
H2O (base model) (sol 1) | 268.10.2 | |||
H2O + clouds (sol 0) | 272.90.2 | |||
H2O + clouds (sol 1) | 269.10.2 | |||
H2O + Stellar activity | 269.60.2 | |||
H2O + Stellar activity + clouds | 271.60.2 |
@@
H2O (base model) | 239.40.2 | |||
H2O + clouds (sol 0) | 227.50.2 | |||
H2O + clouds (sol 1) | 227.40.2 | |||
H2O + Stellar activity | 266.20.2 | |||
H2O + Stellar activity + clouds | 265.60.3 | |||
@@
H2O (base model) | 242.60.2 | |||
H2O + clouds | 242.00.2 | |||
H2O + Stellar activity | 263.80.3 | |||
H2O + Stellar activity + clouds | 263.20.3 |
@@
H2O (base model) | 311.70.2 | |||
H2O + clouds (sol 0) | 312.60.2 | |||
H2O + clouds (sol 1) | 298.60.2 | |||
H2O + Stellar activity | 309.70.3 | |||
H2O + Stellar activity + clouds | 314.10.3 | |||
@@
H2O (base model) | 286.90.2 | |||
H2O + clouds (sol 0) | 289.20.3 | |||
H2O + clouds (sol 1) | 288.60.2 | |||
H2O + Stellar activity | 289.70.3 | |||
H2O + Stellar activity + clouds | 313.80.3 |
@@
H2O (base model) | 304.50.2 | |||
H2O + clouds (sol 0) | 305.90.2 | |||
H2O + clouds (sol 1) | 302.90.2 | |||
H2O + clouds + K (sol 0) | 306.50.2 | |||
H2O + clouds + K (sol 1) | 303.30.2 | |||
H2O + Stellar activity | 311.80.2 | |||
H2O + Stellar activity + clouds | 312.40.2 | |||
H2O + Stellar activity + clouds + K | 312.60.2 | |||
@@
H2O (base model) | 302.90.2 | |||
H2O + clouds (sol 0) | 302.60.2 | |||
H2O + clouds (sol 1) | 302.10.2 | |||
H2O + clouds + K (sol 0) | 304.00.2 | |||
H2O + clouds + K (sol 1) | 302.60.2 | |||
H2O + Stellar activity | 303.50.2 | |||
H2O + Stellar activity + clouds | 304.10.2 | |||
H2O + Stellar activity + clouds + K | 305.20.2 |
@@
H2O (base model) | 289.60.2 | |||
H2O + clouds | 291.00.2 | |||
H2O + Stellar activity | 288.70.3 | |||
H2O + Stellar activity + clouds | 289.70.2 | |||
@@
H2O (base model) | 288.10.2 | |||
H2O + clouds (sol 0) | 288.60.2 | |||
H2O + clouds (sol 1) | 285.60.2 | |||
H2O + Stellar activity | 284.90.3 | |||
H2O + Stellar activity + clouds | 286.30.3 |
@@
H2O (base model) | 286.20.2 | |||
H2O + clouds (sol 0) | 293.10.2 | |||
H2O + clouds (sol 1) | 291.10.2 | |||
H2O + clouds + K (sol 0) | 292.90.2 | |||
H2O + clouds + K (sol 1) | 290.80.2 | |||
H2O + Stellar activity | 295.90.2 | |||
H2O + Stellar activity + clouds | 298.60.2 | |||
H2O + Stellar activity + clouds + K | 298.30.2 | |||
@@
H2O (base model) | 303.00.2 | |||
H2O + clouds (sol 0) | 304.20.2 | |||
H2O + clouds (sol 1) | 302.10.2 | |||
H2O + clouds + K | 303.90.2 | |||
H2O + Stellar activity | 305.70.2 | |||
H2O + Stellar activity + clouds | 308.20.2 | |||
H2O + Stellar activity + clouds + K | 308.00.2 |
@@
H2O (base model) | 293.60.2 | |||
H2O + clouds (sol 0) | 303.60.2 | |||
H2O + clouds (sol 1) | 307.50.2 | |||
H2O + Stellar activity (sol 0) | 297.20.2 | |||
H2O + Stellar activity (sol 1) | 306.30.3 | |||
H2O + Stellar activity + clouds | 307.60.3 | |||
@@
H2O (base model) (sol 0) | 302.80.2 | |||
H2O (base model) (sol 1) | 306.50.2 | |||
H2O + clouds | 314.20.2 | |||
H2O + Stellar activity (sol 0) | 313.00.3 | |||
H2O + Stellar activity (sol 1) | 304.90.3 | |||
H2O + Stellar activity + clouds | 314.20.3 |
@@
H2O (base model) (sol 0) | 255.70.2 | |||
H2O (base model) (sol 1) | 253.40.2 | |||
H2O + clouds | 268.60.2 | |||
H2O + clouds + K + Na | 273.70.3 | |||
H2O + Stellar activity | 255.20.3 | |||
H2O + Stellar activity + clouds | 280.90.3 | |||
H2O + Stellar activity + clouds + Na + K | 281.50.3 |